OpenStudy (anonymous):

write y=(-3/4)x+3 in standard form using integers. Please help:D will fan and medal

jimthompson5910 (jim_thompson5910):

to start things off, you can multiply both sides by the LCD 4 to get y=(-3/4)x+3 4*y=4*[(-3/4)x+3] 4y = 4*(-3/4)*x + 4*3 4y = (-12/4)x + 12 4y = -3x + 12 Notice how the fraction went away

jimthompson5910 (jim_thompson5910):

now you need to get the equation into the form Ax+By = C

OpenStudy (anonymous):

how do i do that?

jimthompson5910 (jim_thompson5910):

A,B,C are just fixed numbers for example, 2x+3y = 7 is in Ax+By = C form where A = 2, B = 3, C = 7

jimthompson5910 (jim_thompson5910):

how would you change 4y = -3x + 12 to get it into the form Ax+By = C? or put another way, how would you change 4y = -3x + 12 so that it looks like 2x+3y = 7

OpenStudy (anonymous):

-3x+4y=12?

jimthompson5910 (jim_thompson5910):

you can't just move the -3x over like that

jimthompson5910 (jim_thompson5910):

you have to do something to both sides to move it over

OpenStudy (anonymous):

What do i have to move? i have so many rules mixed up in my head i cant think of how to do anything rn.

jimthompson5910 (jim_thompson5910):